The minute hand of a clock is 4.5 inches long .what distance does the tip move in 25 minutes


1
Expert's answer
2022-01-14T15:23:47-0500

Solution:


s = r * θ where s is the arc length, r is the radius of the circle and θ is the angle in radians.

s = 4.5in * θ

25 mins are "\frac{5}{12}" of an hour and the angle is "\frac{5}{12}" of "2 \pi = \frac{5 \pi}{6}"


s = 4.5in * "\frac{5 \pi}{6}" = "\frac{15 \pi}{6}" inches, about 11.78 inches